OpenStudy (anonymous):

Find Cos2X, Sin2X, Tan2X. If CosX=-3 and TanX<0

Parth (parthkohli):

\[\cos(2x) = 1- 2\cos^2 x\]

Parth (parthkohli):

Wait, how is \(\cos(x) = -3\)...?

OpenStudy (anonymous):

It was stated in the question. CosX=-3 and tanX is negative

Parth (parthkohli):

As we all know, \(-1 \le \cos x \le 1\). @satellite73 Add another one to the list...

OpenStudy (anonymous):

I assume it just means that this lies in the second quadrant? if cosX is negative and tan is negative.

Parth (parthkohli):

What I mean is that \(\cos(x)\) can never be -3. Never, ever, ever, ever. How can the base be three times the hypotenuse in magnitude? Isn't the hypotenuse supposed to have the biggest magnitude?

OpenStudy (anonymous):

I guess this is a confusing question. its straight from my textbook though. So not sure what to say. I definitely have a 1 for my hypotenuse so far.